OpenStudy (nickymail48):

9-x+5x-3-9x how to solve

OpenStudy (anonymous):

P:arentheses E:xponets M:ultipulcation D:ivision A:ddition S:ubtraction P.E.M.D.A.S

OpenStudy (anonymous):

9-x+5x-3-9x add same terms.. (9-3)+(5-1-9)x =6+(5-10)x =6-5x @keepsecure93

OpenStudy (anonymous):

first combine like terms so 5x,-x, and -9x = -5x and 9-3 which =6 it should be -5x+6

OpenStudy (nickymail48):

thanks

OpenStudy (anonymous):

Hi @nickymail48 . I hope I can help you! The (easiest) way I learned to solve a problem like this is to make a table based on the type of terms. So you have x-terms and constants. Make columns labeled that way and then put the entire term into its respective column. x | c _________ -x | 9 +5x |- 3 -9x | Solve downwards using a calculator. For the 'x' column, just solve the coefficients. So -1 plus 5 minus 9. (then add the 'x' to the sum). = -5x. For the 'c' column, do the same but don’t add anything. 9 minus 3. = 6. Put these two together, making sure to put them in the correct order. -5x + 6. 9 - x + 5x - 3 - 9x = -5x + 6 I hope that’s what you were looking for! :)